# Assembly Theory is an approximation to algorithmic complexity based on LZ compression that does not explain selection or evolution

### v5

@@ A.3.3 Assembly index is a version of LZ compression
  • $t\right>$ is an encoding of the $t$-th member of $\mathbb{P}\left(\mathcal{T}\right)$, where $\mathbb{P}\left(\mathcal{T}\right)$ is the powerset of possibles edges of the minimum directed acyclic multigraph $\mathcal{T}$ into which any possible minimum rooted assembly subspace (whose longest rooted paths have length $\left|\gamma_{y}^{\max}\right|$) can be embedded, and all the most distant vertices from the terminal vertex (i.e., $y$) can only be the vertices in $B_{\Gamma}$.
@@
as input. Then, by employing $\mathbf{F}$, it constructs the directed acyclic multigraph (DAmG) $\mathcal{T}$ of height $\left|\gamma_{y}^{\max}\right|$, and enumerates all the possible minimum rooted assembly subspaces contained in $\mathcal{T}$ as the following procedure: it starts from those that are isomorphic to an arbitrarily fixed $\gamma_{y}^{\max}$ with any possible combination of initial vertex/object in $B_{\Gamma}$; and then in a dovetailing procedure, it indexes first those subspaces that differ (from this $\gamma_{y}^{\max}$) by the least amount of additional edges. Then, program $p_{5}$ chooses the $t$-th first $\Gamma^{\prime}$ that appears in this enumeration. In the next step, program $p_{5}$ reads the rest of the sequence $S_{AT}\left(y\right)$, extracting from the encoded tuples the sequence $I=\left({i_{t}}_{1},\dots,{i_{t}}_{\left|\gamma_{y}^{\min}\right|},{v_{l}}_{1},\dots,{v_{l}}_{\left|\gamma_{y}^{\min}\right|}\right)$ of indexes of edges that are contained in this $t$-th first $\Gamma^{\prime}$ such that it constitutes a shortest rooted path that starts at vertex $v_{B}$. Finally, it uses this shortest rooted path to assemble the final object $y^{\prime}$ accordingly. This describes the decoding procedure from $S_{AT}\left(y\right)$. In order to construct a program $p^{\prime}_{5}$ that encodes a minimum rooted assembly subspace ${\Gamma^{*}}_{y}$ into a $S_{AT}\left(y\right)$, one just needs to note that the codewords for $\left|\gamma_{y}^{\max}\right|\right>$, $B_{\Gamma}\right>$, and $y\right>$ are unique and always retrievable by an algorithm given ${\Gamma^{*}}_{y}$ as input. From these, program $p^{\prime}_{5}$ can use the same procedures that $p_{5}$ ran in order to construct $\mathcal{T}$. Program $p^{\prime}_{5}$ then employs the same procedures that $p_{5}$ ran in order to enumerate the minimum rooted assembly subspaces embedded into $\mathcal{T}$ so as to find the index $t$ that corresponds to ${\Gamma^{*}}_{y}$ in this enumeration. Next, it enumerates all the possible shortest rooted path in this $t$-th minimum rooted assembly subspace. Then, it extracts $v_{B}$ and $I$ from $\mathcal{T}$ by matching those that correspond to the edges and their respective labels in the first shortest rooted path $\gamma_{y}^{\min}$ found in the previous enumeration that terminates at the object $y$. Finally, the program returns the sequence $S_{AT}\left(y\right)$ as output. ∎
@@
With the purpose of demonstrating that $\Gamma^{\prime}={\Gamma^{*}}_{y}$, first note that there is only a finite number of rooted assembly spaces whose most distant vertices (from the vertex/object $y$) belong to $B_{\Gamma}$ (which is also finite and fixed given $\Gamma$) and are separated by $\left|\gamma_{y}^{\max}\right|$ hops. Therefore, since $\left|\gamma_{y}^{\max}\right|\right>$ and $B_{\Gamma}\right>$ are unique from ${\Gamma^{*}}_{y}$ and the enumeration procedure is the same in both the encoding or the deconding phase, there is a unique index $t$ that corresponds to $\Gamma^{\prime}={\Gamma^{*}}_{y}$. In order to achieve the proof of Equation (14), one can exploit the minimality of ${\Gamma^{*}}_{y}$ and $\gamma_{y}^{\min}$ and the maximality of $\gamma_{y}^{\max}$. From any rooted assembly space one of whose longest rooted paths is $\gamma_{y}^{\max}$, in order to construct $\mathcal{T}$, an algorithm can minimise the number of sufficient and necessary new vertices that needs to be added to $\gamma_{y}^{\max}$ by the following procedure: assign an index $j\in\mathbb{N}$ to each of the edges in $\gamma_{y}^{\max}$ by starting from the closest ones to the farthest ones from the initial vertex, which is the object in the basis set $B_{\Gamma}$, such that $1\leq j\leq\left|\gamma_{y}^{\max}\right|$; for the label of the edge $j=1$ in $\gamma_{y}^{\max}$, one only requires at most one extra vertex, and this vertex can only belong to $B_{\Gamma}$ because otherwise, since there would need to be an additional edge connecting this new vertex to the target vertex of edge $j=1$, it would contradict the maximality of $\gamma_{y}^{\max}$; the same analogously holds for each arbitrary $j+1\leq\left|\gamma_{y}^{\max}\right|$, so that labelling this edge $j+1$ with a vertex only requires at most one extra new vertex, and this vertex can only be at level $\leq j$, i.e., it can only belong to the set of vertices that are separated from the terminal vertex $y$ by at least $\left|\gamma_{y}^{\max}\right|-j+1$ hops (otherwise, it would again contradict the maximality of $\gamma_{y}^{\max}$); and for each new vertex created at level $j$ one would only need at most $2$ new vertices at level $\leq j-1$, and so on. Thus, by induction, one will have it that
@@
In accordance with the authors in [1, 41, 53, 2, 8], Theorem 9 reveals that Assembly Theory’s fundamental intuition to classify the notion of complexity of assembled objects is not only that of a compression scheme, but specifically a LZ scheme. The non-unidimensional “tree”-like structure of the minimum rooted assembly subspaces is one of the ways to define a scheme that encodes the assembling process itself back into a linear sequence of codewords.
The encoding-decoding scheme behind Equation (15) translates the notion of the minimum assembling path in an assembly space playing a role in how much simpler the object can get (i.e., as we have demonstrated, how much more compressible the object is). Shorter minimum paths lead to more compressed forms of the assembling process of the objects. The exponential growth in the LZ encoding from the assembly index method is, in turn, consonant with the assembly number [2] of an ensemble (see Section A.3.4).
@@
For example, the reader is invited to note that due to the fact that $\gamma_{y}^{\max}=\gamma_{y}^{\min}$ would hold, a linear minimum rooted assembly subspace ${\Gamma^{*}}_{y}$ (such as in [1, Figure 1(a)] or [53, Figure 8]) results in a much tighter upper bound in Equation (14). This is because the exponential dependence on the assembly index to encode all possible DAG-like structures would drop: that is, $1\leq t\leq\mathbf{O}\left(\left|\gamma_{y}^{\max}\right|^{2}\,\left|B_{\Gamma}\right|\right)$ will hold instead of Equation (19), which in turn will be subsumed in the rightmost term of Equation (14) where one already has a quadratic dependence on $\left|\gamma_{y}^{\max}\right|$ or $c_{\Gamma}\left(y\right)$.
@@
Thus, Theorem 9 not only demonstrates that those minimum rooted assembly subspaces (from which the assembly indexes are calculated) are LZ schemes, but also that the intractability of the calculation of the assembly index (or, equivalently, the intractability of the encoding-decoding procedure) derives from the notion of complexity that the assembly index intends to measure. Together with the height of the grammars in Theorem 7, this also captures the notion of “depth” that AT intends to attain but that in fact pertains, as we have demonstrated, to a compression scheme and a generative grammar. The more complex the DAG-like structure of ${\Gamma^{*}}_{y}$ is, the more distinct DAG-like structures are possible, and therefore the closer the upper bound becomes to the exponential dependence. Hence, Equation (14) would render the object (or assembling process) less compressible, which demonstrates the claims underlying Assembly Theory that the authors in [1, 41, 53, 2, 8] originally intended to grasp with the assembly index but that in fact we demonstrated with a compression scheme. These results show that the notion of how complex an assembling process needs to be in order to construct an object is indeed equivalent to how much less compressible (by a LZ scheme) this process is.
Further, as in Section A.3.4, one can demonstrate that any measure defined upon the assembly index (so that it increases when the assembly index(es) also increases), such as the assembly number [2], also defines an encoding and a compression method.
@@ A.3.4 The assembly number defines an encoding and a compression scheme
Notice that each term $e^{a_{i}}$ plays the role of weights for the respective ratio/proportion $\frac{n_{i}-1}{N_{T}}\leq 1$. In this manner, one can, for example, normalise again for ensembles whose overall summation of assembly indexes remains the same, hence only differing by the concentration/distribution of the quantity of objects with higher assembly index value. As stated in [1, 41, 53, 2], grasping this distinction in fact is the main goal that Assembly Theory aims at with its methods, so as to measure the presence of constraints or biases that turn otherwise unlikely objects into objects that have occurred more often.
For any two ensembles $X$ and $X^{\prime}$ with the same total number $N_{T}$ of objects and whose overall summation $\sum_{i=1}^{N}e^{a_{i}}$ of assembly indexes is also the same, the ensemble with the higher concentration of high-assembly-index objects (with respect to the concentration of these objects in the other ensemble) should score higher on the measure that assembly theory aims to develop. This prevents a (n equal-size) comparison between a biased sampling and another sampling in which the former was taken under distinct conditions in the latter. In this case, without normalising by the summation $\sum_{i=1}^{N}e^{a_{i}}$, one could have the first one scoring a higher assembly number than the second one, but only because of the set of unique objects, and not because of the distribution of higher-assembly-index objects with respect to lower-assembly-index objects, which was what Assembly Theory intended in the first place.777 In order to avoid the presence of false positives due to abiotic mimic [63, 64], an even better control group for excluding sampling condition biases in experiments could be achieved by only comparing ensembles with not only equal size and equal total summation of assembly indexes but also the same set of unique objects. Notice that the forthcoming results in Theorem 10 with regards to assembly number hold analogously for any further normalisation than the one already given by Equation (21). For example, to tackle this issue, the value given by the measure
@@
In case the set of possible ensembles can be arbitrarily large and the ensembles are not generated by stationary and ergodic sources—for example, in this case the minimum expected codeword length per symbol may not converge to the entropy rate—, finding the optimal code is not as straightforward as the one that achieved Equation (22). To tackle this generalised scenario, algorithmic complexity theory improves on any entropic-based coding methods, or on any pure statistical compression method. By using algorithmic information theory (AIT), from which algorithmic complexity is one of its indexes, one obtains the universally optimal encoding, and hence one demonstrates in Theorem 10 and Corollary 11 that the assembly number value is equivalent to the degree of compressibility of the ensemble. That is, the assembly number calculation method is subsumed into the universal coding of AIT, demonstrating that the assembly number in fact captures the idea of compressiblity of the ensemble—not to be conflated with the object—, although only as an approximation that can be arbitrarily distorted in general (see Section A.2.1).
@@
From AIT [21], ones knows that if an arbitrary sequence $\left(\left(d_{1},\tau_{1}\right),\dots,\left(d_{i},\tau_{i}\right),\dots\right)$ of pairs is enumerable and $\sum\limits_{i}2^{-d_{i}}\leq 1$, then one can encode each $\tau_{i}$ in a codeword of length $d_{i}$. Additionally, one also has it that
@@
As proposed by [1, 41, 53, 2, 8], our results demonstrate that ensembles [^10] Not to be conflated with the objects themselves. with higher assembly numbers [^11] Not to be conflated with the assembly indexes. are more compressible, and therefore they diverge more from those ensembles that are outcomes of perfectly random processes. A more random generative process for an ensemble (for which the assembly number scores lower) would mean that there are fewer constraints or biases playing a role in this process. In turn, high-assembly-index objects would more rarely appear (or be constructed) because they are more complex (i.e., as we have demonstrated in the previous sections, less compressible or more incompressible)—this is what occurs in scenarios in which there is a weaker presence of top-down (or downward) causation [11] behind the possibilities or paths that lead to the construction of the objects. As demonstrated in this section, if one assumes that assembly number indeed quantifies the presence of constraints or biases in the underlying generative processes of the ensembles, then a more compressible ensemble implies that more constraints or biases played a role in generating more unlikely (i.e., less compressible) objects more frequently than would have been the case in an environment with otherwise fewer constraints and biases (i.e., a more random or incompressible environment), hence increasing the frequency of occurrence of high-assembly-index (i.e., less compressible) objects in this environment. Conversely, under the same assumption, if more biotic (or less random abiotic) processes being present in an ensemble imply a higher assembly number, then more biotic processes being present in an ensemble would also imply that the ensemble is more compressible.
@@
A.4 Lack of empirical evidence in favour of Assembly Theory and empirical evidence against it
The experiments and analysis of organic versus non-organic molecules from mass spectral data using several complexity indexes used as positive evidence for Assembly Theory in [41] empirically demonstrates what we have proven. That the assembly index is equivalent to statistical compression, given that it was not compared to any other data type or any other existing index of the same nature. The extent of the simplicity of the experiments required to prove the utility of the assembly index and the main hypothesis of Assembly Theory makes us believe that this was a serious omission.
Together with the theoretical results, Figure 3 from [4] shows that the evidence that AT presents as being in favour of their theory and algorithms, specifically that their assembly (compression) index may correspond to the way objects assemble, is flawed and invalid. The empirical evidence actually shows that other, equally plausible and valid compression schemes that are sequential, count copies, and build a repeating symbol dictionary tree, are likely the ones from which AT’s discriminatory power derives and on which AT concepts and methods are based.
Figure 3: Taken from [4]. MA stands for Molecular Assembly based on the assembly index of the molecular compounds. These results were obtained even without making any optimisations to the various algorithms after a trivial flattening procedure of the data (which preserves algorithmic complexity but can help the non-MA indexes converge faster).
The strongest positive correlation in Fig. 3, was identified between MA and 1D-RLE coding (R= 0.9), which is one of the most basic coding schemes known in computer science since the 60s and among the most similar to the intended definition of MA, as being capable of ‘counting copies’ together with LZ that is exactly equivalent to MA or the assembly index. All indexes were applied to spectral data of the 132 molecules (114 plus 18 living extracts) from the original paper in [41] made available by the authors of AT. Other coding algorithms, including the Huffman coding (R = 0.896), also show a strong positive correlation with MA. The analysis further confirms our previous findings of the similarity in performance between MA and other basic compression measures (that essentially only ‘count copies’ in data, all of them converging to Shannon Entropy) in classifying living vs. non-living mass spectra signatures. All molecules included in the Assembly Theory paper were included [41]. Code available in [4].

### v6

@@ A.3.3 The assembly index is a version of LZ compression
  • $t\right>$ is an encoding of the $t$-th member of $\mathbb{P}\left(\mathcal{T}\right)$, where $\mathbb{P}\left(\mathcal{T}\right)$ is the powerset of possible edges of the minimum directed acyclic multigraph $\mathcal{T}$ into which any possible minimum rooted assembly subspace (whose longest rooted paths have length $\left|\gamma_{y}^{\max}\right|$) can be embedded, and all the most distant vertices from the terminal vertex (i.e., $y$) can only be the vertices in $B_{\Gamma}$.
@@
as input. Then, by employing $\mathbf{F}$, it constructs the directed acyclic multigraph (DAmG) $\mathcal{T}$ of height $\left|\gamma_{y}^{\max}\right|$, and enumerates all the possible minimum rooted assembly subspaces contained in $\mathcal{T}$ via the following procedure: it starts from those that are isomorphic to an arbitrarily fixed $\gamma_{y}^{\max}$ with any possible combination of initial vertex/object in $B_{\Gamma}$; and then in a dovetailing procedure, it indexes first those subspaces that differ (from this $\gamma_{y}^{\max}$) by the least amount of additional edges. Then, program $p_{5}$ chooses the $t$-th first $\Gamma^{\prime}$ that appears in this enumeration. In the next step, program $p_{5}$ reads the rest of the sequence $S_{AT}\left(y\right)$, extracting from the encoded tuples the sequence $I=\left({i_{t}}_{1},\dots,{i_{t}}_{\left|\gamma_{y}^{\min}\right|},{v_{l}}_{1},\dots,{v_{l}}_{\left|\gamma_{y}^{\min}\right|}\right)$ of indexes of edges that are contained in this $t$-th first $\Gamma^{\prime}$ such that it constitutes a shortest rooted path that starts at vertex $v_{B}$. Finally, it uses this shortest rooted path to assemble the final object $y^{\prime}$ accordingly. This describes the decoding procedure from $S_{AT}\left(y\right)$. In order to construct a program $p^{\prime}_{5}$ that encodes a minimum rooted assembly subspace ${\Gamma^{*}}_{y}$ into a $S_{AT}\left(y\right)$, one just needs to note that the codewords for $\left|\gamma_{y}^{\max}\right|\right>$, $B_{\Gamma}\right>$, and $y\right>$ are unique and always retrievable by an algorithm given ${\Gamma^{*}}_{y}$ as input. From these, program $p^{\prime}_{5}$ can use the same procedures that $p_{5}$ ran in order to construct $\mathcal{T}$. Program $p^{\prime}_{5}$ then employs the same procedures that $p_{5}$ ran in order to enumerate the minimum rooted assembly subspaces embedded into $\mathcal{T}$ so as to find the index $t$ that corresponds to ${\Gamma^{*}}_{y}$ in this enumeration. Next, it enumerates all the possible shortest rooted paths in this $t$-th minimum rooted assembly subspace. Then, it extracts $v_{B}$ and $I$ from $\mathcal{T}$ by matching those that correspond to the edges and their respective labels in the first shortest rooted path $\gamma_{y}^{\min}$ found in the previous enumeration that terminates at the object $y$. Finally, the program returns the sequence $S_{AT}\left(y\right)$ as output. ∎
@@
With the purpose of demonstrating that $\Gamma^{\prime}={\Gamma^{*}}_{y}$, first note that there is only a finite number of rooted assembly spaces whose most distant vertices (from the vertex/object $y$) belong to $B_{\Gamma}$ (which is also finite and fixed given $\Gamma$) and are separated by $\left|\gamma_{y}^{\max}\right|$ hops. Therefore, since $\left|\gamma_{y}^{\max}\right|\right>$ and $B_{\Gamma}\right>$ are distinct from ${\Gamma^{*}}_{y}$ and the enumeration procedure is the same in both the encoding and the decoding phase, there is a unique index $t$ that corresponds to $\Gamma^{\prime}={\Gamma^{*}}_{y}$. In order to achieve the proof of Equation (14), one can exploit the minimality of ${\Gamma^{*}}_{y}$ and $\gamma_{y}^{\min}$ and the maximality of $\gamma_{y}^{\max}$. From any rooted assembly space one of whose longest rooted paths is $\gamma_{y}^{\max}$, in order to construct $\mathcal{T}$, an algorithm can minimise the number of sufficient and necessary new vertices that need to be added to $\gamma_{y}^{\max}$ using the following procedure: assign an index $j\in\mathbb{N}$ to each of the edges in $\gamma_{y}^{\max}$ starting from the closest and moving to the farthest ones from the initial vertex, which is the object in the basis set $B_{\Gamma}$, such that $1\leq j\leq\left|\gamma_{y}^{\max}\right|$; for the label of the edge $j=1$ in $\gamma_{y}^{\max}$, one only requires at most one extra vertex, and this vertex can only belong to $B_{\Gamma}$ because otherwise, since there would need to be an additional edge connecting this new vertex to the target vertex of edge $j=1$, it would contradict the maximality of $\gamma_{y}^{\max}$; the same analogously holds for each arbitrary $j+1\leq\left|\gamma_{y}^{\max}\right|$, so that labelling this edge $j+1$ with a vertex only requires at most one extra new vertex, and this vertex can only be at level $\leq j$, i.e., it can only belong to the set of vertices that are separated from the terminal vertex $y$ by at least $\left|\gamma_{y}^{\max}\right|-j+1$ hops (otherwise, it would again contradict the maximality of $\gamma_{y}^{\max}$); and for each new vertex created at level $j$ one would only need at most $2$ new vertices at level $\leq j-1$, and so on. Thus, by induction, one will have it that
@@
In relation to the claims made by the authors of Assembly Theory[1, 42, 54, 2, 8], Theorem 9 reveals that Assembly Theory’s fundamental intuition to classify the notion of complexity of assembled objects is not only that of a compression scheme, but specifically an LZ scheme. The non-unidimensional “tree”-like structure of the minimum rooted assembly subspaces is one of the ways to define a scheme that encodes the assembling process itself back into a linear sequence of codewords.
The encoding-decoding scheme behind Equation (15) translates the notion of the minimum assembling path in an assembly space, playing a role in how much simpler the object can get (i.e., as we have demonstrated, how much more compressible the object is). Shorter minimum paths lead to more compressed forms of the assembling process of the objects. The exponential growth in the LZ encoding from the assembly index method is, in turn, consonant with the assembly number [2] of an ensemble (see Section A.3.4).
@@
For example, the reader is invited to note that due to the fact that $\gamma_{y}^{\max}=\gamma_{y}^{\min}$ would hold, a linear minimum rooted assembly subspace ${\Gamma^{*}}_{y}$ (such as in [1, Figure 1(a)] or [54, Figure 8]) results in a much tighter upper bound in Equation (14). This is because the exponential dependence on the assembly index to encode all possible DAG-like structures would drop: that is, $1\leq t\leq\mathbf{O}\left(\left|\gamma_{y}^{\max}\right|^{2}\,\left|B_{\Gamma}\right|\right)$ will hold instead of Equation (19), which in turn will be subsumed in the rightmost term of Equation (14) where one already has a quadratic dependence on $\left|\gamma_{y}^{\max}\right|$ or $c_{\Gamma}\left(y\right)$.
@@
Thus, Theorem 9 not only demonstrates that those minimum rooted assembly subspaces (from which the assembly indexes are calculated) are LZ schemes, but also that the intractability of the calculation of the assembly index (or, equivalently, the intractability of the encoding-decoding procedure) derives from the notion of complexity that the assembly index intends to measure. Together with the height of the grammars in Theorem 7, this also captures the notion of “depth” that AT intends to attain but that in fact pertains, as we have demonstrated, to a compression scheme and a generative grammar. The more complex the DAG-like structure of ${\Gamma^{*}}_{y}$ is, the more distinct DAG-like structures are possible, and therefore the closer the upper bound becomes to the exponential dependence. Hence, Equation (14) would render the object (or assembling process) less compressible, which demonstrates the claims underlying Assembly Theory that the authors in [1, 42, 54, 2, 8] originally intended to demonstrate with the assembly index but that in fact we demonstrated with a compression scheme. These results show that the notion of how complex an assembling process needs to be in order to construct an object is indeed equivalent to how much less compressible (by a LZ scheme) this process is.
Further, as in Section A.3.4, one can demonstrate that any measure defined in terms of the assembly index (so that it increases when the assembly index(es) also increases), such as the assembly number [2], also defines an encoding and a compression method. The higher the assembly value of an object, the more the object is LZ compressible and the other way around, in direct and equal proportion.
@@ A.3.4 The assembly number defines an encoding and a compression scheme
Notice that each term $e^{a_{i}}$ plays the role of weight for the respective ratio/proportion $\frac{n_{i}-1}{N_{T}}\leq 1$. In this manner, one can, for example, normalise again for ensembles whose overall summation of assembly indexes remains the same, hence only differing by the concentration/distribution of the quantity of objects with higher assembly index value. As stated in [1, 42, 54, 2], grasping this distinction in fact is the main goal that Assembly Theory aims at with its methods, so as to measure the presence of constraints or biases that turn otherwise unlikely objects into objects that occur more often.
For any two ensembles $X$ and $X^{\prime}$ with the same total number $N_{T}$ of objects and whose overall summation $\sum_{i=1}^{N}e^{a_{i}}$ of assembly indexes is also the same, the ensemble with the higher concentration of high-assembly-index objects (with respect to the concentration of these objects in the other ensemble) should score higher on the measure that assembly theory aims to develop. This prevents a (n equal-size) comparison between a biased sampling and another sampling in which the former was taken under distinct conditions in the latter. In this case, without normalising by the summation $\sum_{i=1}^{N}e^{a_{i}}$, one could have the first one scoring a higher assembly number than the second one, but only because of the set of unique objects, and not because of the distribution of higher-assembly-index objects with respect to lower-assembly-index objects, which was what Assembly Theory intended in the first place. [^7] In order to avoid the presence of false positives due to abiotic mimicry [64, 65], an even better control group for excluding sampling condition biases in experiments could be achieved by only comparing ensembles with not only equal size and equal total summation of assembly indexes but also the same set of unique objects. Notice that the forthcoming results in Theorem 10 with regards to assembly number hold analogously for any further normalisation than the one already given by Equation (21). For example, to tackle this issue, the value given by the measure
@@
In case the set of possible ensembles can be arbitrarily large and the ensembles are not generated by stationary and ergodic sources—for example, in this case the minimum expected codeword length per symbol may not converge to the entropy rate—finding the optimal code is not as straightforward as finding the code that generated Equation (22). To tackle this generalised scenario, algorithmic complexity theory improves on any entropy-based coding methods, or on any pure statistical compression method. By using algorithmic information theory (AIT), algorithmic complexity being one of its indexes, one obtains the universally optimal encoding, and hence one demonstrates in Theorem 10 and Corollary 11 that the assembly number value is equivalent to the degree of compressibility of the ensemble. That is, the assembly number calculation method is subsumed into the universal coding of AIT, demonstrating that the assembly number, in fact, captures the idea of compressiblity of the ensemble—not to be conflated with the object— although only as an approximation that can, in general, be arbitrarily distorted (see Section A.2.1).
@@
From AIT [21], one knows that if an arbitrary sequence $\left(\left(d_{1},\tau_{1}\right),\dots,\left(d_{i},\tau_{i}\right),\dots\right)$ of pairs is enumerable and $\sum\limits_{i}2^{-d_{i}}\leq 1$, then one can encode each $\tau_{i}$ in a codeword of length $d_{i}$. Additionally, one also has it that
@@
As proposed by [1, 42, 54, 2, 8], our results demonstrate that ensembles [^10] Not to be conflated with the objects themselves. with higher assembly numbers [^11] Not to be conflated with the assembly indexes. are more compressible, and therefore they diverge more from those ensembles that are outcomes of perfectly random processes. A more random generative process for an ensemble (for which the assembly number scores lower) would mean that there are fewer constraints or biases playing a role in this process. In turn, high-assembly-index objects would more rarely appear (or be constructed) because they are more complex (i.e., as we have demonstrated in the previous sections, less compressible or more incompressible)—this is what occurs in scenarios in which there is a weaker presence of top-down (or downward) causation [11] behind the possibilities or paths that lead to the construction of the objects. As demonstrated in this section, if one assumes that the assembly number indeed quantifies the presence of constraints or biases in the underlying generative processes of the ensembles, then a more compressible ensemble implies that more constraints or biases played a role in generating more unlikely (i.e., less compressible) objects more frequently than would have been the case in an environment with fewer constraints and biases (i.e., a more random or incompressible environment), hence increasing the frequency of occurrence of high-assembly-index (i.e., less compressible) objects in this environment. Conversely, under the same assumption, if more biotic (or less random abiotic) processes being present in an ensemble imply a higher assembly number, then more biotic processes being present in an ensemble would also imply that the ensemble is more compressible.
@@
A.4 Lack of empirical evidence in favour of Assembly Theory and presence of empirical evidence against it
The experiments and analysis of organic versus non-organic molecules from mass spectral data using several complexity indexes used as positive evidence for Assembly Theory in [42] empirically demonstrates what we have proven: that the assembly index is equivalent to statistical compression, given that it was not compared to any other data type or any other existing index of the same nature. The simplicity of the experiments required to prove the utility of the assembly index and the main hypothesis of Assembly Theory inclines us to believe that this was a serious omission.
Together with the theoretical results, Figure 3 from [4] shows that the evidence that the AT authors present as being in favour of their theory and algorithms, specifically that their assembly (compression) index may correspond to the way objects assemble, is flawed and invalid. The empirical evidence actually shows that other, equally plausible and valid compression schemes that are sequential, count copies, and build a repeating symbol dictionary tree, are likely the ones from which AT’s discriminatory power derives and on which AT’s concepts and methods are based.
Figure 3: Taken from [4]. The most basic statistical complexity indexes applied to the 18 molecular compounds for which the spectral data was made available from [42]. MA stands for Molecular Assembly based on the assembly index of the molecular compounds. These results were obtained even without optimising the algorithms after a simple flattening and bnarisation procedure of the data (full methods, data and code available at [4]). The results show that other statistical measures separate the data just as MA does (or better) between living and non-living compounds.
The strongest positive correlation in Fig. 3, was identified between MA and 1D-RLE coding (R= 0.9), which is one of the most basic coding schemes in computer science, known since the 60s, and among the most similar to the intended definition of MA, as being capable of ‘counting copies’, together with LZ, which is exactly equivalent to MA or the assembly index. All indexes were applied to spectral data (only 18 extracts available in the original paper in [42]) and taking their results at face value. Other coding algorithms, including the Huffman coding (R = 0.896), also show a strong positive correlation with MA. The analysis further confirms our previous findings of the similarity in performance between MA and other basic compression measures (that essentially only ‘count copies’ in data, all of them converging to Shannon Entropy) in classifying living vs. non-living mass spectra signatures. All molecules included in the Assembly Theory paper were included. In other words, both theory and experiment converge, and the experiments conform with the findings in this paper that Assembly Theory is equivalent to Shannon Entropy and their assembly index is an LZ compression algorithm renamed.